OpenStudy (anonymous):

What is the equation for the axis of symmetry in the function? y = -2x2 + 5x - 4 A. y = 5/4 B. x = 5/4 C. y = -2/5 D. x = -2/5

sam (.sam.):

Try completing the square

OpenStudy (anonymous):

would it be x=5/4?

sam (.sam.):

\[-2x^2 + 5x - 4\] \[-2(x^2 - \frac{5x}{2}) - 4\] \[-2(x - \frac{5}{4})^2-(\frac{25}{16})(-2) - 4\] \[-2(x - \frac{5}{4})^2-\frac{7}{8}\] symmetry at x= 5/4

sam (.sam.):

yes

OpenStudy (anonymous):

thanks!

sam (.sam.):

welcome :)